# SOOD++: Leveraging Unlabeled Data to Boost Oriented Object Detection

### v1

@@ 3.1 Pseudo-labeling Paradigm
In this paper, we consider the limitations of both sparse and dense pseudo-label sampling strategies when dealing with small-scale, densely distributed remote-sensing objects. Specifically, the SPL may fail to detect small objects, resulting in insufficient supervisory information. Conversely, DPL, despite its abundance, may introduce substantial noise due to the illegibility between dense small objects and the background. Thus, we propose a Simple Instance-aware Dense Sampling (SIDS) strategy while sampling dense pseudo-labels from the predicted boxes of foreground and background, generating more comprehensive dense pseudo-labels.
@@ 3.2 Optimal Transport
where $\mathbf{1}_{n}$ is a $n$-dimensional column vector with all elements equal to 1. The constraints ($\boldsymbol{P}\mathbf{1}_{n}=\hat{\boldsymbol{x}}$ and $\boldsymbol{P}^{T}\mathbf{1}_{n}=\hat{\boldsymbol{y}}$) ensure the probability mass from each point in the source distribution $\hat{\boldsymbol{x}}$ is accurately transported to the target distribution $\hat{\boldsymbol{y}}$, maintaining the total mass invariant. The Monge-Kantorovich’s Optimal Transport (OT) cost between $\hat{\boldsymbol{x}}$ and $\hat{\boldsymbol{y}}$ is then defined as:
@@
[…]
@@ 4 Method
Fig. 2 illustrates the pipeline of our SOOD++. Specifically, the SOOD++ is a dense pseudo-labeling framework, which consists of three key components: a Simple Instance-aware Dense Sampling (SIDS) strategy is used to construct the high-quality dense pseudo-labels, the Geometry-aware Adaptive Weighting (GAW) loss and Noise-driven Global Consistency (NGC) are adopted to measure the discrepancy between the teacher and student from a one-to-one and set-to-set (global) perspectives, respectively. In the following contents, we describe the overall framework in Sec. 4.1, and introduce the key design of the proposed components, i.e., SIDS, GAW, and NGC, in the following Sec. 4.2, Sec. 4.3, and Sec. 4.4, respectively.
Figure 2: The pipeline of our SOOD++. Each training batch includes labeled and unlabeled images, omitting the regular supervised part for simplicity. For the unsupervised part, we use the Simple Instance-aware Dense Sampling (SIDS) strategy to create high-quality pseudo-labels and then pair them with the student’s predictions. The Geometry-aware Adaptive Weighting (GAW) loss dynamically weighs each pair’s unsupervised loss based on geometry information. Additionally, we treat the pseudo-labels and student predictions as global discrete distributions, measuring their similarity via Noise-driven Global Consistency (NGC). Yellow points indicate sampled dense pseudo-labels and predictions.
@@ 4.1 The Overall Framework
Since aerial objects are usually dense and small, the sparse-pseudo-labeling paradigm might miss massive potential objects. Thus, we choose the dense pseudo-labeling (DPL) paradigm. The training process includes the supervised and unsupervised parts. For the supervised part, the student is trained regularly with labeled data. For the unsupervised part, we adopt the following steps:
@@
We adopt the widely-used rotated version of FCOS [56] as the teacher and student. Note that at each pixel-level point, rotated-FCOS will predict the confidence score, centerness, and bounding box. The basic unsupervised loss contains classification loss, regression loss, and center-ness loss, corresponding to the output of FCOS. We adopt smooth $\mathcal{L}_{1}$ loss for regression loss, Binary Cross Entropy (BCE) loss for classification and center-ness losses. Based on these, our GAW and NGC will be used to measure the consistency of the teacher and student from different perspectives.
@@ 4.2 Simple Instance-aware Dense Sampling Strategy
Figure 3: (a) GT box. (b) Sparse pseudo-label (bounding box) [64, 40]. (c) Previous methods [76, 35] sample dense pseudo-labels from the entire score map using Top-K. (d) We sample dense pseudo-labels from the predicted instance (foreground) and background with dynamic numbers.
@@
Previous dense pseudo-labeling methods [76, 35] typically extract a fixed number (i.e., Top-K) of dense pseudo-labels from the predicted score map (Fig. 3(c)). In contrast, we dynamically sample a variable number of dense pseudo-labels from the predicted instance (box) areas and background (Fig. 3(d)). Specifically, we categorize the pseudo-labels from unlabeled data into two types: the easy case obtained from the predicted results and the hard case, which often blends seamlessly into the background. Given the predicted results from the teacher after post-processing (e.g., NMS), considered as the foreground (easy case), we randomly sample pixel-level prediction results within the region indicated by these predicted boxes with a sample ratio $\mathcal{R}$, resulting in $\mathcal{P}^{e}$. Note that due to the dynamic nature of the number of instances, our sampling number varies across different scenes, better matching the distribution of the objects.
@@
However, merely adopting the above strategy to sample dense pseudo-labels from the foreground often overlooks dense, small objects that blend seamlessly into the background. To address this issue, we propose an intuitive approach to extract potential dense pseudo-labels from the background. Specifically, we empirically find that those potential high-quality positive samples, despite their low confidence, align well with the target location. Consequently, we can use the Intersection over Union (IoU) score between the predicted box and ground-truth box as a criterion for discerning pseudo-labels amidst background noise. However, given the lack of ground truth in the unlabeled dataset, explicit IoU scores are unattainable. Therefore, we introduce an IoU estimation branch trained with the labeled data that predicts the IoU of each pixel for the unlabeled data. Then, we apply threshold $\mathcal{T}_{h}$ to choose the pixel-level points with high IoU scores from the background, resulting in dense hard pseudo-labels $\mathcal{P}^{h}$. We merge $\mathcal{P}^{h}$ and $\mathcal{P}^{e}$ to establish the final dense pseudo-labels $\mathcal{P}^{t}$, which will serve as the supervisory signal for the unsupervised training process. Based on $\mathcal{P}^{t}$, we sample the predictions from the student at the same coordinates, defined as $\mathcal{P}^{s}$. After obtaining $\mathcal{P}^{t}$ and $\mathcal{P}^{s}$, the semi-supervised framework aims to make $\mathcal{P}^{s}$ approach $\mathcal{P}^{t}$, and we will describe how to construct one-to-one and many-to-many constraints in the following sections.
@@
Admittedly, IoU estimation is indeed a widely used technique. However, our objective fundamentally differs from previous methods [13, 26, 30], which primarily focus on box refinement. In contrast, our method aims to mine potential dense pseudo-labels from the background in a semi-supervised framework. It should be noted that this paper does not claim the novelty of the proposed SIDS but rather presents a simple and effective sampling strategy that yields high-quality dense pseudo-labels.
@@ 4.3 Geometry-aware Adaptive Weighting Loss
Orientation is one essential property of an oriented object. As shown in Fig. 1(a), their orientations remain clear even if objects are dense and small. Some oriented object detection methods [46, 68, 72] have already employed such property in loss calculation. These works are under the assumption that the angles of the labels are reliable. Strictly forcing the prediction close to the ground truth is natural.
@@
Based on the above observations, we propose to softly utilize orientation information and aspect ratios, enabling the semi-supervised framework to better understand the geometric priors of oriented objects. Specifically, The difference in orientation between a prediction and a pseudo-label can indicate the difficulty of the instance to some extent and can be used to adjust the unsupervised loss dynamically. Moreover, since oriented objects with different aspect ratios exhibit varying sensitivities to rotation changes, this can be leveraged to modulate the importance of orientation differences in indicating instance difficulty. Therefore, we propose a geometry-aware modulating factor based on these two critical geometric priors. Similar to focal loss [34], this factor dynamically weights the loss of each pseudo-label-prediction pair by considering its intrinsic geometric information (e.g., orientation difference and aspect ratio).
@@
where $r_{i}^{t}$ and $r_{i}^{s}$ represent the rotation angles in radians of the $i$-th pseudo-label and prediction, respectively. $a_{i}^{t}$ and $a_{i}^{s}$ denote their aspect ratios. Adding a constant to $\sigma_{i}$ ensures that the original unsupervised loss is preserved when the orientations of the pseudo-label and prediction coincide. The modulating factor $\sigma_{i}$ considers both orientation differences and average aspect ratios. A toy example is that when the difference in orientation is marginal, but the object’s aspect ratio is large, the proposed factor can still efficiently highlight such a case as a possible hard example and help improve the learning process, and vice versa.
@@
where $N_{p}$ is the number of pseudo-labels and $\mathcal{L}_{u}^{i}$ is the basic unsupervised loss of the $i$-th pseudo-label-prediction pair. Using the geometry-aware modulating factor, the GAW loss provides a clear direction for optimization, facilitating the semi-supervised learning process.
Figure 4: IoU vs. Rotation angle over different aspect-ratios.
@@ 4.4 Noise-driven Global Consistency
The objects in an aerial image are usually dense and regularly distributed, as shown in Fig. 1(a). We argue that such an arrangement is similar to the structured layout of the text within a document. The spatial configuration of the object set is implicitly referred to as the layout, conveying the inter-object relationships and the overarching pattern intrinsic to the image. Ideally, the layout consistency between the student’s and the teacher’s predictions will be preserved if each pseudo-label-prediction pair is aligned. However, such a condition is too strict and may hurt performance when there are noises in pseudo-labels. Therefore, it is reasonable to add the consistency between layouts as an additional relaxed optimization objective, encouraging the student to learn information from the teacher from a global perspective. Besides, by treating the predictions from students as a global distribution, the relations between these predictions can be regularized implicitly, which provides implicit constraint to the student to some extent.
@@
The details of our Noise-driven Global Consistency (NGC) are shown in Fig. 5. Given the output of teacher and student, we will treat them as two global distributions. First, we add different random noises to disturb both the teacher and the student distributions. Then, we propose to make multi-perspective alignments from three aspects: 1) aligning distributions between the original teacher and original student; 2) aligning distributions between the disturbed teacher and disturbed student; 3) aligning the transport plans from 1) and 2). Applying the multi-perspective alignments is beneficial for maintaining consistency and stability during the learning process. In other words, this global constraint makes the output distribution of the student model as close as possible to that of the teacher model despite disturbance. It will increase the models’ tolerance to minor variations in inputs, i.e., allowing them to more effectively capture the essential features of the unlabeled data without depending on noise or specific details.
Specifically, we propose to leverage the optimal transport theory to measure the global similarity of layouts between the teacher’s and the student’s predictions. Specifically, let $K$ denote the number of classes. We define the global distributions of classification scores predicted by the teacher ($\mathbf{s}^{t}\in\mathbb{R}^{N_{p}\times K}$) and the student ($\mathbf{s}^{s}\in\mathbb{R}^{N_{p}\times K}$) as follows:
@@
where $cls(i)\in\{0,1,\cdots,K-1\}$ is the class index with the largest score for the $i$-th pseudo-label. And $\mathbf{s}_{i,cls(i)}^{t}$ indicates the score of $cls(i)$-th class for the $i$-th pseudo-label. The $\mathbf{s}_{i,cls(i)}^{s}$ has a similar definition but refers to the predictions of the student. The exponential used here is empirical, aiming for numerical stability and ensuring the value is greater than 0.
@@
where ($\boldsymbol{\lambda}^{*},\boldsymbol{\mu}^{*})$ and ($\widetilde{\boldsymbol{\lambda}}^{*},\widetilde{\boldsymbol{\mu}}^{*}$) are the solutions of Eq. 3. We solve the OT problem by a fast Sinkhorn distances algorithm [9], obtaining the approximate solution. Based on the defined $\mathcal{L}_{GC}(\mathbf{d}^{t},\mathbf{d}^{s})$ and $\mathcal{L}_{GC}(\widetilde{\mathbf{d}}^{t},\widetilde{\mathbf{d}}^{s})$, their gradients [^2] Note that under the semi-supervised learning framework, only the predictions of the student model have gradients. with respect to $\mathbf{d}^{s}$ and $\widetilde{\mathbf{d}}^{s}$ follows:
@@
$\mathcal{L}_{plan}=\mathrm{MSE}(\boldsymbol{P}-\widetilde{\boldsymbol{P}})$ (16)
@@
when we optimize such OT-based loss, the predictions converge closely to the pseudo-labels from a global perspective. We argue that promoting such multiple global consistency not only relieves the negative effect of inaccuracy pseudo-labels when detection processing, but also encourages a holistic learning approach, considering objects’ spatial and relational dynamics.

### v2

@@ 3.2 Optimal Transport
where $\mathbf{1}_{n}$ is a $n$-dimensional column vector with all elements equal to 1. The constraints ($\boldsymbol{P}\mathbf{1}_{n}=\hat{\boldsymbol{x}}$ and $\boldsymbol{P}^{T}\mathbf{1}_{n}=\hat{\boldsymbol{y}}$) ensure the probability mass from each point in the source distribution $\hat{\boldsymbol{x}}$ is accurately transported to the target distribution $\hat{\boldsymbol{y}}$, maintaining the total mass invariant. The cost between $\hat{\boldsymbol{x}}$ and $\hat{\boldsymbol{y}}$ is then defined as:
@@
where $\boldsymbol{C}\in\mathbb{R}^{n\times n}$ represents the cost matrix between two sets, and $\left\langle\cdot\right\rangle$ represents inner product. In common, the OT problem is solved by the following dual formulation:
@@ 4 Method
Fig. 3 illustrates the pipeline of our SOOD++. Specifically, the SOOD++ is a dense pseudo-labeling framework, which consists of three key components: a Simple Instance-aware Dense Sampling (SIDS) strategy is used to construct the high-quality dense pseudo-labels, the Geometry-aware Adaptive Weighting (GAW) loss and Noise-driven Global Consistency (NGC) are adopted to measure the discrepancy between the teacher and student from an individual-level and global-level perspectives, respectively.
Figure 3: The pipeline of our SOOD++. Each training batch includes labeled and unlabeled images, omitting the regular supervised part for simplicity. For the unsupervised part, we use the Simple Instance-aware Dense Sampling (SIDS) strategy to create high-quality pseudo-labels and pair them with the student’s predictions. The Geometry-aware Adaptive Weighting (GAW) loss dynamically weighs each pair’s unsupervised loss based on geometry information. Additionally, we treat the pseudo-labels and student predictions as global discrete distributions, measuring their similarity via Noise-driven Global Consistency (NGC). Yellow points indicate sampled dense pseudo-labels and predictions.
@@ 4.1 The Overall Framework
Aerial objects are usually dense and small, whereas the sparse-pseudo-labeling paradigm might miss massive potential objects. Thus, we choose the dense pseudo-labeling paradigm. The training process includes the supervised and unsupervised parts. For the supervised part, the student is trained regularly with labeled data. For the unsupervised part, we adopt the following steps:
@@
We adopt the widely-used rotated version of FCOS [56] as the teacher and student. At each pixel-level point, rotated-FCOS will predict the confidence score, centerness, and bounding box. The basic unsupervised loss contains classification loss, regression loss, and centerness loss, corresponding to the output of FCOS. We adopt smooth $\mathcal{L}_{1}$ loss for regression loss, Binary Cross Entropy (BCE) loss for classification and centerness losses. Based on these, our GAW and NGC will be used to measure the consistency of the teacher and student from different perspectives.
@@ 4.2 Simple Instance-aware Dense Sampling Strategy
Previous dense pseudo-labeling methods [3, 5] typically extract a fixed number (i.e., Top-K) of dense pseudo-labels from the predicted score map (Fig. 2(c)). In the conference version (SOOD) [53], we propose to sample dense pseudo-labels from the foreground box areas (Fig. 2(d)), where the foreground represents the predicted results from the teacher model after post-processing (e.g., NMS). However, it has a limitation: it often overlooks dense, small objects that blend seamlessly into the background, leading to insufficient supervision. For instance, potential high-quality positive samples might align well with the target location but are overlooked due to their low confidence in the predicted score maps.
@@
To address the above limitation, we introduce the Simple Instance-aware Dense Sampling (SIDS) strategy, which categorizes pseudo-labels into easy (from the foreground) and hard (from the background) cases and processes them accordingly (Fig. 2(e)). Specifically, in the foreground, we randomly sample pixel-level prediction results within the region indicated by these predicted boxes with a sample ratio $\mathcal{R}$, resulting in $\mathcal{P}^{e}$. In the background, we find that those potential high-quality positive samples align well with the target location despite their low confidence in the score maps. This inspires us to use the Intersection over Union (IoU) score between the predicted box and ground-truth box as a criterion for discerning pseudo-labels amidst background noise. However, given the lack of ground truth in the unlabeled dataset, explicit IoU scores are unattainable. Besides, the default rotated-FCOS output consists of class scores and bounding box coordinates (i.e., center, width, and height) for each pixel, but it does not directly capture the IoU between the predicted and ground truth bounding boxes. Therefore, we introduce an IoU estimation branch (only one $\times3$ convolutional layer) trained with the labeled data that predicts the IoU of each pixel for the unlabeled data. Then, we apply threshold $\mathcal{T}_{h}$ to choose the pixel-level points with high IoU scores from the background, resulting in dense hard pseudo-labels $\mathcal{P}^{h}$.
We merge $\mathcal{P}^{h}$ and $\mathcal{P}^{e}$ to establish the final dense pseudo-labels $\mathcal{P}^{t}$, which will serve as the supervisory signal for the unsupervised training process. Based on $\mathcal{P}^{t}$, we sample the predictions from the student at the same coordinates, defined as $\mathcal{P}^{s}$. After obtaining $\mathcal{P}^{t}$ and $\mathcal{P}^{s}$, the semi-supervised framework aims to make $\mathcal{P}^{s}$ approach $\mathcal{P}^{t}$, and we will describe how to construct individual-level and global-level constraints in the following sections. Due to the varying number and size of objects, our sampling number varies across different scenes, better matching the distribution of the objects.
@@
Admittedly, IoU estimation is indeed a widely used technique. However, our objective fundamentally differs from previous methods [57, 58, 59], which primarily focus on box refinement. In contrast, our method aims to mine potential dense pseudo-labels from the background in a semi-supervised framework.
@@ 4.3 Geometry-aware Adaptive Weighting Loss
The orientation of the oriented objects is a distinct and easily identifiable characteristic, even in cases where the objects are dense and small, as shown in Fig. 1(a). Some oriented object detection methods [60, 61, 62] have already employed such property in loss calculation. These works are under the assumption that the angles of the labels are reliable. Strictly forcing the prediction close to the ground truth is natural.
@@
Based on the above observations, we propose to softly utilize orientation information and aspect ratios, enabling the semi-supervised framework to better understand the geometric priors of oriented objects. Specifically, The difference in orientation between a prediction and a pseudo-label can indicate the difficulty of the instance to some extent, which can be used to adjust the unsupervised loss dynamically. Moreover, considering that oriented objects with different aspect ratios exhibit varying sensitivities to rotation changes, it inspires us to use the aspect ratio to modulate the importance of orientation differences in indicating instance difficulty. Therefore, we propose a geometry-aware modulating factor based on these two critical geometric priors. Similar to focal loss [27], this factor dynamically weights the loss of each pseudo-label-prediction pair by considering its intrinsic geometric information (e.g., orientation difference and aspect ratio).
Figure 4: IoU vs. Rotation angle over different aspect-ratios.
@@
where $r_{i}^{t}$ and $r_{i}^{s}$ represent the rotation angles in radians of the $i$-th pseudo-label and prediction, respectively. $a_{i}^{t}$ and $a_{i}^{s}$ denote their aspect ratios. $\psi$ is a loss balance weight. Adding a constant to $\sigma_{i}$ ensures that the original unsupervised loss is preserved when the orientations of the pseudo-label and prediction coincide. The modulating factor $\sigma_{i}$ considers both orientation differences and average aspect ratios. A toy example is that when the difference in orientation is marginal, but the object’s aspect ratio is large, the proposed factor can still efficiently highlight such a case as a possible hard example and help improve the learning process, and vice versa.
@@
where $N_{p}$ is the number of pseudo-labels and $\mathcal{L}_{u}^{i}$ is the basic unsupervised loss of the $i$-th pseudo-label-prediction pair.
Note that our GAW builds upon the Rotation-aware Adaptive Weighting (RAW) of the conference version [53] by incorporating the aspect ratio as an additional factor in the weighting scheme, providing a clearer direction for optimization during the semi-supervised learning.
@@ 4.4 Noise-driven Global Consistency
In aerial scenes, objects often exhibit spatial relationships that are not random. For instance, vehicles might align along roads, or ships may be arranged in ports. The spatial configuration of the object set is implicitly referred to as the layout, conveying the inter-object relationships and the overarching pattern intrinsic to the image. Ideally, the layout consistency between the student’s and the teacher’s predictions will be preserved if each pseudo-label-prediction pair is aligned. However, such a condition is too strict and may hurt performance when there are noises in pseudo-labels. Therefore, it is reasonable to add the consistency between layouts as an additional relaxed optimization objective, encouraging the student to learn information from the teacher from a global perspective. Besides, by treating the predictions from students as a global distribution, the relations between these predictions can be regularized implicitly, which provides an implicit constraint to the student to some extent.
@@
The details of our Noise-driven Global Consistency (NGC) are shown in Fig. 5. Given the output of the teacher and student, we will treat them as two global distributions. First, we add different random noises to disturb both the teacher and the student distributions. Then, we propose to make multi-perspective alignments from three aspects: 1) aligning distributions between the original teacher and original student; 2) aligning distributions between the disturbed teacher and disturbed student; 3) aligning the transport plans from 1) and 2). By adding noise to both the teacher’s and student’s global distribution, we encourage the model to focus on global patterns rather than overfitting to noise or minor inconsistencies in the pseudo-labels. We argue that this regularization improves generalization, especially when the pseudo-labels are imperfect. The multi-perspective alignment ensures consistency at a different global-level perspective, maintaining the overall object structure and preventing the model from relying too heavily on potentially noisy or incorrect pseudo-labels.
Specifically, we propose to leverage the optimal transport theory to measure the global similarity of layouts between the teacher’s and the student’s predictions. Specifically, let $K$ denote the number of classes. For a given image [^3] The student model samples pixel-level results at the same positions as the teacher, ensuring consistency in both the number and location., we define the global distributions of classification scores predicted by the teacher ($\mathbf{s}^{t}\in\mathbb{R}^{N_{p}\times K}$) and the student ($\mathbf{s}^{s}\in\mathbb{R}^{N_{p}\times K}$) as follows:
@@
where $cls(i)\in\{0,1,\cdots,K-1\}$ is the class index with the largest score for the $i$-th pseudo-label. And $\mathbf{s}_{i,cls(i)}^{t}$ indicates the score of $cls(i)$-th class for the $i$-th pseudo-label. The $\mathbf{s}_{i,cls(i)}^{s}$ has a similar definition but refers to the predictions of the student. The exponential used here aims for numerical stability and ensuring the value is greater than 0.
@@
where ($\boldsymbol{\lambda}^{*},\boldsymbol{\mu}^{*})$ and ($\widetilde{\boldsymbol{\lambda}}^{*},\widetilde{\boldsymbol{\mu}}^{*}$) are the solutions of Eq. 3. We solve the OT problem by a fast Sinkhorn distance algorithm [55], obtaining the approximate solution. Based on the defined $\mathcal{L}_{GC}(\mathbf{d}^{t},\mathbf{d}^{s})$ and $\mathcal{L}_{GC}(\widetilde{\mathbf{d}}^{t},\widetilde{\mathbf{d}}^{s})$, their gradients [^4] Note that under the semi-supervised learning framework, only the predictions of the student model have gradients. with respect to $\mathbf{d}^{s}$ and $\widetilde{\mathbf{d}}^{s}$ follows:
@@
[…]
@@
when we optimize such OT-based loss, the predictions converge closely to the pseudo-labels from a global perspective.
@@
In the conference version (SOOD) [53], we have already introduced the concept of global consistency but only considering $\mathcal{L}_{GC}(\mathbf{d}^{t},\mathbf{d}^{s})$. SOOD++ further refines it by introducing random noise perturbations to both the teacher and student distributions. Then, we make multi-perspective global alignments, i.e., including $\mathcal{L}_{GC}(\mathbf{d}^{t},\mathbf{d}^{s})$, $\mathcal{L}_{GC}(\widetilde{\mathbf{d}}^{t},\widetilde{\mathbf{d}}^{s})$, and $\mathcal{L}_{plan}$, as illustrated in Fig. 5. This noise-driven, multi-perspective consistency mechanism ensures that the model is not overly sensitive to local pseudo-label noise. Instead, it encourages learning of robust global features that reflect the overall spatial and relational structure of the scene. By aligning both the original and noisy distributions, the model effectively captures essential global patterns.